OpenStudy (anonymous):

Need some help please.. Obtain the general solution: (D^2 +3D +2)y = 12x^2

OpenStudy (anonymous):

Auxiliary equation is D^2+3D+2=0 D^2+2D+D+2=0 D(D+2)+1(D+2)=0 (D+2)(D+1)=0 D=-1,-2 \[General~ solution =C1e ^{-x}+C2e ^{-2x}\]

OpenStudy (anonymous):

Right, I can get that far but the total answer is: \[y=C1e^{-x} + C2e^{-2x} + 6x^2 -18x +2\] I need help on the 2nd part of the answer please..

OpenStudy (anonymous):

to find particular integral let\[ y=ax^2+bx+c\] \[D=\frac{ dy }{ dx }=2ax+b\] \[D^2=\frac{ d^2D }{ dx^2 }=2a\]

OpenStudy (anonymous):

put in the give eq.and find a,b,c i am leaving.

OpenStudy (anonymous):

So is it.. \[2A +3(2Ax+B)+2(Ax^2+Bx+C)=12x^2\] ?

OpenStudy (anonymous):

The answer i gave, the last number should be a 21

OpenStudy (anonymous):

\[2a+6ax+3b+2(ax^2+bx+c)=12x^2\] \[2ax^2+(2b+6a)x+2a+3b+2c=12x^2\] 2a=12 a=6 2b+6*6=0 \[b=\frac{ -36 }{2 }=-18\] 2a+3b+2c=0 2*6+3*-18+2c=0 2c=54-12=42 \[c=\frac{ 42 }{2 }=21\] \[hence P.I.=6x^2-18x+21\] \[c.s.~ is~y=C1e ^{-x}+C2e ^{-2x}+6x^2-18x+21\]

OpenStudy (anonymous):

You are awesome! Thank you!

OpenStudy (anonymous):

YW